# Exploring the Impact of HAPS-RIS on UAV-Based Networks: a Novel Network Architecture

### v2

@@ III-A HAPS-RIS Zone
The solution of OP 1-b, i.e., $R^{*}$, is obtained in the proposed Algorithm 1 which is an iterative algorithm. In each iteration, the zone boundary and user sets are updated, and users in the HAPS-RIS zone are assigned to their respective RIS clusters using a random injective function, with an equal number of subcarriers allocated to each user. In the proposed geometric channel model for the HAPS-RIS zone, the channel gain is affected by the locations of the RIS clusters, the users, and the CS. Hence, the random allocation of RIS clusters is justified due to the geometric configuration of the system, where the RIS dimensions are negligible compared to the distances between the HAPS and the users. Furthermore, an injective function is adopted to ensure that each user in the HAPS-RIS zone is uniquely assigned to one of the available RIS clusters. This process continues until the maximum number of users in the HAPS-RIS zone is achieved while satisfying the rate requirement.
Algorithm 1 Solving OP 1-b: Obtaining the Zone Boundary
s1:  Initialize $t=0$, $\delta=50$, $R^{*}=R_{0}$, and define the set $\mathcal{I}$={$i|\sqrt{x^{2}_{i}+y^{2}_{i}}\leqslant R_{0}$}
@@
s3:  Define the set ${\mathcal{B}^{(t)}}$={$i^{\prime}|\sqrt{x^{2}_{i^{\prime}}+y^{2}_{i^{\prime}}}\leqslant R^{(t)}$} where $R^{(t)}=R_{0}-t\delta$. $T$ is the upper limit of the iteration count.
@@ III-B UAV Zone
$r^{\rm{UAV}}_{i}=\sum\limits^{J}_{j=1}\sum\limits^{L^{\rm{UAV}}}_{l^{\prime}=1}B_{l}\log_{2}\left(1+\gamma^{l^{\prime},\rm{UAV}}_{ij}\right),\forall i\in\mathcal{B},$ (11)
@@
where the available subcarriers are $L^{\rm{UAV}}$=$\frac{L^{\rm{tot}}}{2}$. The Signal to interference plus noise ratio (SINR) of the user $i$ served by UAV $j$ over subcarrier $l^{\prime}$ can be obtained as
@@
$\begin{array}[]{l}\gamma^{l^{\prime},\rm{UAV}}_{ij}=\frac{P^{l^{\prime},\rm{UAV}}_{ij}\tilde{\rho}^{l^{\prime},\rm{UAV}}_{ij}\left|h^{l^{\prime},\rm{UAV}}_{ij}\right|^{2}}{N_{0}B_{l}+\sum^{J}_{j^{\prime}\neq j}P^{l^{\prime},\rm{UAV}}_{ij^{\prime}}\tilde{\rho}^{l^{\prime},\rm{UAV}}_{ij^{\prime}}\left|h^{l^{\prime},\rm{UAV}}_{ij^{\prime}}\right|^{2}},\forall i\in\mathcal{B},\forall j,l^{\prime}\end{array}$ (12)
@@
where $P^{l^{\prime},\rm{UAV}}_{ij}$ is the transmit power of the UAV $j$ allocated to the subcarrier $l^{\prime}$ and the user $i$. $h^{l^{\prime},\rm{UAV}}_{ij}$ is the channel gain between UAV $j$ and user $i$ over subcarrier $l^{\prime}$, which can be formulated as $h^{\rm{UAV}}_{ij}=\sqrt{G^{\rm{UAV}}_{j}G_{i}\bar{\mathcal{L}}^{-1}_{ij}},\forall i\in\mathcal{B},\forall j$. $G^{\rm{UAV}}_{j}$ is the antenna gain of the UAV $j$, $G_{i}$ is the antenna gain of the user $i$, and ¯$\bar{\mathcal{L}}_{ij}$ is the average path loss which can be obtained as [15]
@@
¯$\begin{array}[]{l}\bar{\mathcal{L}}_{ij}=P^{\rm{UAV}}_{ij,\rm{LoS}}\eta_{1}\left(\frac{4\pi f_{c}}{c}d_{ij}\right)^{\alpha}+P^{\rm{UAV}}_{ij,\rm{NLoS}}\eta_{2}\left(\frac{4\pi f_{c}}{c}d_{ij}\right)^{\alpha},\\ \forall i\in\mathcal{B},\forall j,\end{array}$ (13)
@@
where $\alpha$ is the path loss exponent, $\eta_{1}$ and $\eta_{2}$ are the excessive path losses corresponding to the LoS and non-LoS connections, respectively, $d_{ij}=\sqrt{\left(x_{i}-x^{\rm{}\rm{UAV}}_{j}\right)^{2}+\left(y_{i}-y^{\rm{UAV}}_{j}\right)^{2}+(z^{\rm{UAV}}_{j})^{2}},\forall i\in\mathcal{B},\forall j$ is the distance between the user $i$ and the UAV $j$, and $P^{\rm{UAV}}_{ij,\rm{}\rm{LoS}}$ is the LoS probability between the user $i$ and the UAV $j$ as
$P^{\rm{UAV}}_{ij,\rm{}\rm{LoS}}=\frac{1}{1+\psi e^{-\beta\left(\tilde{\theta}_{ij}-\psi\right)}},\forall i\in\mathcal{B},\forall j,$ (14)
@@
where $\beta$ and $\psi$ are the constant values that depend on the environment type. $P^{\rm{UAV}}_{ij,\rm{}\rm{LoS}}+P^{\rm{UAV}}_{ij,\rm{}\rm{NLoS}}=1,\forall i\in\mathcal{B},\forall j,$ and $\tilde{\theta}_{ij}=\frac{180}{\pi}\arcsin\left(\frac{z^{\rm{UAV}}_{j}}{d_{ij}}\right),\forall i\in\mathcal{B},\forall j$. Considering all these factors, the follower optimization problem can be designed as
$\textbf{OP 2}:\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \min_{N^{\rm{UAV}}}\ N^{\rm{UAV}}$ (15a)
@@
$\begin{array}[]{l}{{\rm{s}}.{\rm{t}}.:}\leavevmode\nobreak\ \leavevmode\nobreak\ r^{\rm{UAV}}_{i}\left(N^{\text{UAV}}\right)\geq r^{\rm{user}}_{0},\ \forall i\in\mathcal{B}.\end{array}$ (15b)
@@
The solution of OP 2, i.e., minimum number of required UAVs, is achieved by using the proposed Algorithm 2 for the given zone boundary, obtained in OP 1. We propose a dynamic method adopting $k$-means clustering technique with adaptive UAV selection. Unlike conventional $k$-means clustering, where the $k$-factor is a fixed input [16], we consider it as a varying number of UAVs that adapts iteratively. This approach increases the network degrees of freedom. In each iteration, the users are associated to the UAVs using $k$-means clustering method while satisfying the subcarrier orthogonality and the rate requirement. The process continues until minimimum number of UAVs is obtained.
Algorithm 2 Solving OP 2: Proposed Dynamic $k$-means Clustering Technique with Adaptive UAV Selection
@@
s1:  Initialize $\delta^{\prime}$=1, $N^{\rm{UAV},(0)}$ and define the set ${\mathcal{B^{*}}}$={$i^{\prime}|\sqrt{x^{2}_{i^{\prime}}+y^{2}_{i^{\prime}}}\leqslant R^{*}$} where $R^{*}$ is the final solution of OP 1.
@@
s4:  Check whether orthogonality of subcarriers are satisfied with considering the constraint $\sum\limits^{I}_{i}\tilde{\rho}^{l^{\prime}}_{ij}\leq 1\,\forall l^{\prime},j$. Then calculate $\boldsymbol{\tilde{\rho}}\left(N^{\text{UAV},(t)}\right)$.
s5: While    $r^{\rm{UAV}}_{i}\left(\boldsymbol{\tilde{\rho}}\left(N^{\text{UAV},(t)}\right),\boldsymbol{x}\left(N^{\text{UAV},(t)}\right),\boldsymbol{y}\left(N^{\text{UAV},(t)}\right)\right)\\ \geq r^{\rm{user}}_{0},\ \forall i\in\mathcal{B}$ or $t<T$, set $N^{*}=N^{{\rm{UAV}},(t)}$.
@@
IV Simulation Results
@@
Fig. 2, Fig. 3, and Fig. 4 are the results which only obtained by solving the leader problem OP 1 when there is only HAPS-RIS to support the users. Fig. 2 and Fig. 3 are the results where a portion of the coverage area is covered by HAPS-RIS and the rest of the users are in the outage. Fig. 4 demonstrates the result where full coverage is provided only by HAPS-RIS system. Fig. 5 is the result obtained by solving both the leader and follower problems, i.e., OP 1 and OP 2, where full coverage is provided by integration of two types of infrastructures, i.e., UAVs and HAPS-RIS system. We take into account a wide range of rate requirements, spanning from kbps to Mbps.
@@
Figure 2: Percentage of covered users by HAPS-RIS versus number of RIS element for the rate requirements 64, 128, and 256 kbps when the CS transmit power is 40 dBm where only HAPS-RIS is exploited.
@@
Figure 3: Percentage of covered users by HAPS-RIS versus CS Transmit power for the rate requirements 64, 128, and 256 kbps when the number of RIS elements are 3.5×3.5$\times$3.5 ×$10^{5}$ where only HAPS-RIS is exploited.
@@
Fig. 2 shows the percentage of users covered by HAPS-RIS versus the number of RIS elements for three different data rate thresholds of 64 kbps, 128 kbps, and 256 kbps. As the number of RIS elements increases, the percentage of covered users also rises across all data rate thresholds. For the lowest data rate of 64 kbps, user coverage begins at approximately 50% and approaches 90% as the number of RIS elements increases. For the 128 kbps threshold, coverage starts at around 20% and increases to about 60%. The highest data rate of 256 kbps shows the lowest coverage, beginning slightly above 10 % and reaching around 30%. Fig. 3 demonstrate the percentage of users covered by HAPS-RIS versus the CS transmit power for various minimum data rate requirements, data rate thresholds of 64 kbps, 128 kbps, and 256 kbps, utilizing $3.5\times 10^{5}$ RIS elements. The findings show that increasing CS transmit power leads to a higher percentage of covered users. These results underscore the pivotal role of both CS transmit power and the number of RIS elements in the coverage provided by HAPS-RIS. Fig. 4 shows that as the minimum rate requirement $r^{\rm{user}}_{0}$ increases, the number of RIS elements $M$ needed for full coverage also rises, both on a logarithmic scale. This indicates that higher rate demands require more RIS elements. The result highlights the potential of the HAPS-RIS system to provide full coverage while satisfying different rate requirements without relying on UAVs.
@@
Fig. 5 presents the minimum number of required UAVs $N^{*}$ versus the total nember of RIS elements $M$ to achieve the full coverage at different minimum rate requirements $r^{\rm{user}}_{0}$. The graph depicts three distinct scenarios: $r^{\rm{user}}_{0}=2$ Mbps, $r^{\rm{user}}_{0}=4$ Mbps, and $r^{\rm{user}}_{0}=8$ Mbps. The stepped nature of each line indicates that as the number of RIS elements increases, the number of UAVs needed decreases in discrete steps. For the highest rate requirement, 8 Mbps, the system requires up to 4 UAVs when the number of RIS elements is low; however, as $M$ increases beyond approximately 5 million elements, no UAVs are needed. Conversely, for lower rate requirements, 2 Mbps and 4 Mbps, fewer UAVs are required overall, and the need for UAVs is eliminated at a lower threshold of RIS elements. This analysis illustrates the trade-off between the number of RIS elements and UAVs: as the number of RIS elements increases, the dependence on UAVs decreases, which is particularly beneficial for meeting higher rate requirements with fewer UAVs.
@@
Figure 4: Required number of total RIS elements in Logarithmic value versus logarithm of the minimum rate requirements to provide full coverage where only HAPS-RIS is exploited.
Figure 5: Minimum number of required UAVs versus total number of RIS elements to provide the full coverage where UAVs and HAPS-RIS is exploited.

### v3

@@ III-A Proposed HAPS-RIS Zone Design
The solution of OP 1-b, i.e., $R^{*}$, is obtained by the proposed DyRaZARC technique. In each iteration, the zone boundary and user sets are updated, and users in the HAPS-RIS zone are assigned to their respective RIS clusters using a random injective function, with an equal number of subcarriers allocated to each user. In the proposed geometric channel model for the HAPS-RIS zone, the channel gain is affected by the locations of the RIS clusters, the users, and the CS. Hence, the random allocation of RIS clusters is justified due to the geometric configuration of the system, where the RIS dimensions are negligible compared to the distances between the HAPS and the users. Furthermore, an injective function is adopted to ensure that each user in the HAPS-RIS zone is uniquely assigned to one of the available RIS clusters. This process continues until the maximum number of users in the HAPS-RIS zone is achieved while satisfying the rate requirement.
The Proposed DyRaZARC Technique: Solving OP 1-b to Obtain the Zone Boundary
s1: Initialize $t=0$, $\delta=50$, $R^{*}=R_{0}$, and define the set $\mathcal{I}$$i|\sqrt{x^{2}_{i}+y^{2}_{i}}\leqslant R_{0}$}
@@
s3: Define the set $\mathcal{B}^{(t)}$$i^{\prime}|\sqrt{x^{2}_{i^{\prime}}+y^{2}_{i^{\prime}}}\leqslant R^{(t)}$} where $R^{(t)}=R_{0}-t\delta$. $T$ is the upper limit of the iteration count.
@@ III-B Proposed UAV Zone Design
$r^{\rm{UAV}}_{i}=\sum\limits^{J=N^{\rm{UAV}}}_{j=1}\sum\limits^{L^{\rm{UAV}}}_{l^{\prime}=1}B_{l}\log_{2}\left(1+\gamma^{l^{\prime},\rm{UAV}}_{ij}\right),\forall i\in\mathcal{B},$ (12)
@@
where the available subcarriers are $L^{\rm{UAV}}$$\frac{L^{\rm{tot}}}{2}$. Hence, the signal to interference plus noise ratio (SINR) of the user $i$ served by UAV $j$ over subcarrier $l^{\prime}$, $\forall i\in\mathcal{B},\forall j,l^{\prime}$, can be obtained as
@@
$\begin{array}[]{l}\gamma^{l^{\prime},\rm{UAV}}_{ij}=\frac{P^{l^{\prime},\rm{UAV}}_{ij}\tilde{\rho}^{l^{\prime},\rm{UAV}}_{ij}\left|h^{l^{\prime},\rm{UAV}}_{ij}\right|^{2}}{N_{0}B_{l}+\sum^{J}_{j^{\prime}\neq j}P^{l^{\prime},\rm{UAV}}_{ij^{\prime}}\tilde{\rho}^{l^{\prime},\rm{UAV}}_{ij^{\prime}}\left|h^{l^{\prime},\rm{UAV}}_{ij^{\prime}}\right|^{2}},\end{array}$ (13)
@@
where $P^{l^{\prime},\rm{UAV}}_{ij}$ is the transmit power of the UAV $j$ allocated to the subcarrier $l^{\prime}$ and the user $i$. $h^{l^{\prime},\rm{UAV}}_{ij}$ is the channel gain between UAV $j$ and user $i$ over subcarrier $l^{\prime}$, which can be formulated as $h^{\rm{UAV}}_{ij}=\sqrt{G^{\rm{UAV}}_{j}G_{i}\bar{\mathcal{L}}^{-1}_{ij}},\forall i\in\mathcal{B},\forall j$. $G^{\rm{UAV}}_{j}$ is the antenna gain of the UAV $j$, $G_{i}$ is the antenna gain of the user $i$, and $\bar{\mathcal{L}}_{ij}$ is the average path loss, $\forall i\in\mathcal{B},\forall j$, which can be obtained as [14]
@@
$\begin{array}[]{l}\bar{\mathcal{L}}_{ij}=\\ P^{\rm{UAV}}_{ij,\rm{LoS}}\eta_{1}\left(\frac{4\pi f_{c}}{c}d_{ij}\right)^{\alpha}+P^{\rm{UAV}}_{ij,\rm{NLoS}}\eta_{2}\left(\frac{4\pi f_{c}}{c}d_{ij}\right)^{\alpha},\par\end{array}$ (14)
@@
where $\alpha$ is the path loss exponent, $\eta_{1}$ and $\eta_{2}$ are the excessive path losses corresponding to the LoS and non-LoS connections, respectively. The distance between the user $i$ and the UAV $j$, $\forall i\in\mathcal{B},\forall j$, can be formulated as
$d_{ij}=\sqrt{\left(x_{i}-x^{\rm{}\rm{UAV}}_{j}\right)^{2}+\left(y_{i}-y^{\rm{UAV}}_{j}\right)^{2}+(z^{\rm{UAV}}_{j})^{2}},$ (15)
and $P^{\rm{UAV}}_{ij,\rm{}\rm{LoS}}$ is the LoS probability between the user $i$ and the UAV $j$ as
$P^{\rm{UAV}}_{ij,\rm{}\rm{LoS}}=\frac{1}{1+\psi e^{-\beta\left(\tilde{\theta}_{ij}-\psi\right)}},$ (16)
@@
where $\beta$ and $\psi$ are the constant values that depend on the environment type. $P^{\rm{UAV}}_{ij,\rm{}\rm{LoS}}+P^{\rm{UAV}}_{ij,\rm{}\rm{NLoS}}=1,\forall i\in\mathcal{B},\forall j,$ and $\tilde{\theta}_{ij}=\frac{180}{\pi}\arcsin(\frac{z^{\rm{UAV}}_{j}}{d_{ij}}),\forall i\in\mathcal{B},\forall j$.
Remark 2: In the UAV zone, we exploit orthogonal frequency division multiple access (OFDMA), and hence, there are multi UAV cells where there is inter-cell interference but no intra-cell interference.
Definition 2: $\boldsymbol{\tilde{\rho}}$ is obtained based on $N^{\rm{UAV}}$ via $k$-means clustering and satisfying subcarrier orthogonality in each UAV cluster. Considering all these factors, the follower optimization problem can be designed as
$\hskip-113.81102pt\textbf{OP 2}:\penalty 10000\ \penalty 10000\ \penalty 10000\ \min_{N^{\rm{UAV}}}\ N^{\rm{UAV}}$ (17a)
@@
$\begin{array}[]{l}{{\rm{s}}.{\rm{t}}.:}\penalty 10000\ \penalty 10000\ r^{\rm{UAV}}_{i}\left(N^{\text{UAV}}\right)\geq r^{\rm{user}}_{0},\ \forall i\in\mathcal{B}.\end{array}$ (17b)
@@
The solution of OP 2, i.e., minimum number of required UAVs, is achieved by exploiting the proposed geometry-informed machine learning KADUS technique, for the given zone boundary which is obtained in OP 1. Unlike conventional $k$-means clustering, where the $k$-factor is a fixed input [15], the proposed KADUS technique considers it as a varying number of UAVs that adapts iteratively. This approach increases the network degrees of freedom. In each iteration, the users are associated to the UAVs using $k$-means clustering method while satisfying the subcarrier orthogonality and the rate requirement. The process continues until minimum number of UAVs is obtained.
The Proposed Geometry-informed Machine Learning KADUS Technique: Solving OP 2
@@
s1: Initialize $\delta^{\prime}$=1, $N^{\rm{UAV},(0)}$ and define the set $\mathcal{B^{*}}$$i^{\prime}|\sqrt{x^{2}_{i^{\prime}}+y^{2}_{i^{\prime}}}\leqslant R^{*}$} where $R^{*}$ is the final solution of OP 1-b.
@@
s4: Apply orthogonality check for subcarriers based on the constraint $\sum\limits^{I}_{i}\tilde{\rho}^{l^{\prime}}_{ij}\leq 1\,\forall l^{\prime},j$. Then calculate $\boldsymbol{\tilde{\rho}}\left(N^{\text{UAV},(t)}\right)$.
s5: While $t<T$ or $r^{\rm{UAV}}_{i}\left(\boldsymbol{\tilde{\rho}}\left(N^{\text{UAV},(t)}\right),\boldsymbol{x}\left(N^{\text{UAV},(t)}\right),\boldsymbol{y}\left(N^{\text{UAV},(t)}\right)\right)\geq r^{\rm{user}}_{0},\ \forall i\in\mathcal{B}$ , set $N^{*}=N^{{\rm{UAV}},(t)}$.
@@
Corollary 1: The solution of the proposed hierarchical bilevel optimization framework including the leader problem, OP 1 which is defined in Definition 1, and the follower one, OP 2 which is defined in Definition 2, is $(R,\boldsymbol{\vartheta},N^{\text{UAV}})=(R^{*},\boldsymbol{\vartheta^{*}},N^{*})$. $\boldsymbol{\vartheta^{*}}$ can be obtained via Proposition 1 where the RIS phase shifts are practically designed in OP 1-a. With the given $\boldsymbol{\vartheta^{*}}$, $R^{*}$ can be obtained via Proposition 2 and the proposed DyRaZARC technique. Then, with the given $R^{*}$, $N^{*}$ can be obtained by the proposed geometry-informed machine learning KADUS technique.
IV Numerical Evaluations
@@
Fig. 2, Fig. 3, and Fig. 4 are results obtained by solving the leader problem OP 1 by exploiting the proposed DyRaZARC technique when there is only HAPS-RIS to support the users. Fig. 2 and Fig. 3 are the results where a portion of the coverage area is covered by HAPS-RIS and the rest of the users are in the outage. Fig. 4 demonstrates the result where full coverage is provided only by HAPS-RIS system. Fig. 5 is the result obtained by exploiting the proposed DyRaZARC and KADUS techniques to solve the leader and follower problems, i.e., OP 1 and OP 2, respectively. In this case, full coverage is provided by integration of two types of infrastructures, i.e., UAVs and HAPS-RIS system. We take into account a wide range of rate requirements, spanning from kbps to Mbps.
@@
Figure 2: Percentage of covered users by HAPS-RIS versus number of RIS element for the rate requirements 64, 128, and 256 kbps when the CS transmit power is 40 dBm and only HAPS-RIS setup is in use.
@@
Figure 3: Percentage of covered users by HAPS-RIS versus CS Transmit power for the rate requirements 64, 128, and 256 kbps when the number of RIS elements are $3.5\times$$10^{5}$ and only HAPS-RIS setup is in use.
IV-A Only HAPS-RIS is in Use
@@
Fig. 2 demonstrates the percentage of users covered by HAPS-RIS versus the number of RIS elements for three different data rate thresholds of 64 kbps, 128 kbps, and 256 kbps. As the number of RIS elements increases, the percentage of covered users also rises across all data rate thresholds. For the lowest data rate of 64 kbps, user coverage begins at approximately 50% and approaches 90% as the number of RIS elements increases. For the 128 kbps threshold, coverage starts at around 20% and increases to about 60%. The highest data rate of 256 kbps shows the lowest coverage, beginning slightly above 10% and reaching around 30%. Fig. 3 demonstrate the percentage of users covered by HAPS-RIS versus the CS transmit power for various minimum data rate requirements, data rate thresholds of 64 kbps, 128 kbps, and 256 kbps, utilizing $3.5\times 10^{5}$ RIS elements. The findings show that increasing CS transmit power leads to a higher percentage of covered users. These results underscore the pivotal role of both CS transmit power and the number of RIS elements in the coverage provided by HAPS-RIS. Fig. 4 shows that as the minimum rate requirement $r^{\rm{user}}_{0}$ increases, the number of RIS elements $M$ needed for full coverage also rises, both on a logarithmic scale. This indicates that higher rate demands require more RIS elements. The result highlights the potential of the HAPS-RIS system to provide full coverage while satisfying different rate requirements without relying on UAVs.
IV-B HAPS-RIS and UAVs are in Use
@@
Fig. 5 presents the minimum number of required UAVs $N^{*}$ versus the total number of RIS elements $M$ to achieve the full coverage at different minimum rate requirements $r^{\rm{user}}_{0}$. The graph depicts three distinct scenarios: $r^{\rm{user}}_{0}=2$ Mbps, $r^{\rm{user}}_{0}=4$ Mbps, and $r^{\rm{user}}_{0}=8$ Mbps. The stepped nature of each line indicates that as the number of RIS elements increases, the number of UAVs needed decreases in discrete steps. For the highest rate requirement, 8 Mbps, the system requires up to 4 UAVs when the number of RIS elements is low; however, as $M$ increases beyond approximately 5 million elements, no UAVs are needed. Conversely, for lower rate requirements, 2 Mbps and 4 Mbps, fewer UAVs are required overall, and the need for UAVs is eliminated at a lower threshold of RIS elements. This analysis illustrates the trade-off between the number of RIS elements and UAVs: as the number of RIS elements increases, the dependence on UAVs decreases, which is particularly beneficial for meeting higher rate requirements with fewer UAVs.
@@
Figure 4: Required number of total RIS elements in logarithmic scale versus logarithm of the minimum rate requirements to provide full coverage, where only HAPS-RIS setup is in use.
Figure 5: Trade-off between the number of UAVs and RIS elements to provide the full coverage. This is achieved by exploiting the proposed DyRaZARC and KADUS techniques where both HAPS-RIS setup and UAVs are in use.